Calculate the amount of concentrated 11.6 M HCl that is necessary to prepare 500.00 mL of a 1.00 M HCl solution? Report your answer to the correct number of significant figures.
Answer #1

Before dilution After dilution M = IM M = 11-6 M V = ? y = 500 ml My = M, L 11-6 X Y = 1.000 500.00 1.00 x 500.00 v = volume